[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jul 15, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 15 10:24:03 UTC 2016


HSU Shift Log for Fri Jul 15 2016 06:23:57 GMT-0400 (EDT)

Shift: A

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None. No obvious problems with the a-decks this morning.

N-AWIPS/AWIPS2 issues:

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160715/2b376a40/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list